Status Report

ISS On-Orbit Status 21 Aug 2002

By SpaceRef Editor
August 21, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted previously or below.

Workday again ran through (ground) night, starting at 4:00 pm EDT yesterday with morning inspection, breakfast and DPC (Daily Planning Conference), and ending at 7:30 am this morning. This sleep cycle is in anticipation of the next Orlan spacewalk, EVA-8, on 8/26, to begin at 1:00 am EDT, as dictated by Russian ground site comm coverage constraints.

CDR Valery Korzun and FE-1 Peggy Whitson then started out in the Service Module (SM) on the newly introduced periodic inspection of the TVIS (treadmill with vibration isolation and stabilization) chassis and its components, going by a 12-page list of procedural instructions uplinked yesterday, and including a brief inspection of the chassis truss surface for a suspected crack. [The two-hour inspection involved the removal of forward and rear deck assemblies, verifying running belt tension, then untensioning it and adjusting the aft drum to be able to lift the belt up and inspect the orientation of the rivet-nuts fasteners of its slats underneath, inspecting the front and rear drum set screws, the truss rollers and the chassis interior, retensioning the belt, installing the forward and rear deck assemblies, and cleaning up.]

Assisted by a ground specialist via S-band telecon, the two cosmonauts, Korzun and Treschev, started their preparations for next Monday’s modified EVA by reviewing procedures and the preliminary timeline, which incorporates the changes caused by the curtailment of last Friday’s spacewalk.

Later, Valery and Sergey completed a training session of the Russian crew health procedure of Hand Ergometry (MO-6), a countermeasure to the decrease in arm muscle tone noticed previously by Russian physical fitness experts in cosmonauts. The tolerance hand grip training/analysis has become a standard Russian pre-EVA requirement, which needed to be included in today’s onboard schedule in time for the 8/26 spacewalk..

Peggy Whitson meanwhile performed monthly maintenance on the CEVIS (cycle ergometer with vibration isolation and stabilization), which essentially consists of examining the wire rope isolators for damage. She found one frayed and one broken cable on both the right forward and right aft end of the CEVIS and, in awaiting further instructions, temporarily protected the frayed wires with gray tape.

She also was scheduled to prepare two Nikon F5 cameras for the EVA, one with a 35 and the other with a 28mm lens.

FE-2 Sergey Treschev performed the periodic maintenance test of closing the Vozdukh CO2 scrubber’s emergency vacuum valves (AVK).

Both Korzun and Treschev completed their weekly data collection session for the joint NTXN Interactions ("Vzaimodejstviye") experiment, which collects personal mood, group and journal impressions for psychological evaluation after the return of the crew.

The daily routine tasks of SM life support systems (SOSh) maintenance, Lab payload status checkup and IMS (inventory management system) delta file preparation were today assigned to Peggy Whitson.

All crewmembers performed their regular physical exercise on TVIS, RED (expander), VELO (cycle ergometer with load trainer) and CEVIS.

CSA (Canadian Space Agency) has reported that the SSRMS/Canadarm2 contractor, MD-R in Brampton, has completed the failure investigation of the SSRMS Wrist Roll Joint brought back on UF-2. The Failure Analysis Report is currently being finalized and will be briefed out to CSA/NASA on 8/28.

Update on Progress propellants: 8P still contains 710 kg of ISS-transferable fuel/oxidizer, of the original amount of 824 kg.

The US CEO (crew earth observations) program today had the following targets: Jakarta, Indonesia (major Third World city right of track on the N coast of densely populated island of Java. Crew was asked to photograph urban margins and look for possible smog plume extending out over the sea, then try for fires in western Borneo [left and right of track] which remains unusually clear), Kabul, Afghanistan (nadir pass), Baghdad, Iraq (nadir pass; of interest: city margins and numerous irrigation canals), Lake Poopo, Bolivia (this very full lake is on the cusp of starting a multi-year, El Nino-related contraction [just right of track]), Lower Amazon River Basin (detailed views of coastlines and islands needed), Western Mediterranean Dust (approaching front was setting up dusty winds from the Sahara. Of interest: plumes with sea as backdrop), Danube valley floods (massive floods in central Europe should be visible–look along major rivers for muddy water flooding croplands, railroads, etc.)

CEO images can be viewed at the website

U.S. and Russian Segment Status (as of 5:15 am EDT):

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):
Elektron O2 generator is powered On (32-amp mode), on backup pump. Vozdukh CO2 scrubber is ON in MANUAL cycle mode #5, i.e., 10-min. cycle time (vacuum pump failed). U.S. CDRA CO2 scrubber is on Override. BMP Harmful Impurities unit: Absorbent bed #1 in Purify mode, bed #2 in Purify mode.

Electrical Power Systems (EPS):
Both P6 channels fully operational. Beta Gimbal Assembly (BGA) 2B and BGA 4B in Autotrack mode (solar-tracking).
SM batteries: Battery #1 is off-line; all other batteries (7) are in "Partial Charge" mode.
FGB batteries: Battery #5 is off-line, battery #2 is in "Cycle" mode; all other batteries (4) are in "Partial Charge" mode.
Plasma Contactor Unit PCU-1 in Discharge mode; PCU-2 in Discharge mode.

Thermal Control Systems:
Air conditioner SKV-1 is On. SKV-2 is Off.

Command & Data Handling Systems:
C&C-3 MDM is prime, C&C-2 is back-up, and C&C-1 is in standby.
GNC-1 MDM is prime; GNC-2 is Backup.
LA-1, LA-2 and LA-3 MDMs are all operating.
PL-1 MDM is operational; PL-2 MDM is Off.
APS-1 (automated payload switch #1) and APS-2 are both On.
SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Attitude Source:
3 CMGs on-line.
State vector — US GPS (SIGI string 1)
Attitude — Russian segment
Angular rates — US RGA1 (rate gyro assembly 1)

Communications & Tracking Systems:
All Russian communications & tracking systems are nominal.
S-band is operating nominally.
Ku-band is operating nominally.
Audio subsystem operating nominally.
Video subsystem operating nominally.
MCOR (medium-rate communications outage recorder) is operating nominally.

SSRMS/Canadarm2 at MBS PDGF 1 (mobile base system/power & data grapple fixture 1), in EVA-7 viewing position, with Keep Alive power on both strings (based on MBS).
MBS: Keep Alive power on both strings.
RWS (robotics workstations): Lab RWS is Off; Cupola RWS is Off.

ISS Orbit (as of this morning, 4:11am EDT [= epoch]):
Mean altitude — 392.9 km
Apogee — 404.5 km
Perigee — 381.3 km
Period — 92.4 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.0017093
Orbits per 24-hr. day — 15.58
Altitude decrease — 250 m (mean) in last 24 hours
Solar Beta Angle — -35.0 deg (magnitude increasing)
Revolutions since FGB/Zarya launch (Nov. ’98) — 21418
Current Flight Attitude — XPOP (x-axis perpendicular to orbit plane = "sun-fixed" [yaw: -0.6 deg, pitch: -5.7 deg., roll: 0 deg]).

For more on ISS orbit and worldwide naked-eye visibility dates/times, see

SpaceRef staff editor.